# Text-Guided Multi-Property Molecular Optimization with a Diffusion Language Model

### v1

@@ 3.2. Description Generator
Molecular Structure Feature.   Molecular structure feature effectively enhances insights into molecules and model performance in specific tasks  (Ferreira et al., 2015; Wang et al., 2024; Djeddi et al., 2023). Among them, fragment-based molecular structure feature has garnered substantial attention from researchers  (Green et al., 2021; Yang et al., 2021). Accordingly, we disassemble molecules using Recap  (Lewell et al., 1998), which cleaves the structures into fragments based on retrosynthetic chemistry. Once key fragments are recognized as essential for chemical reactivity, they can easily be divided into latent blocks for optimization. After molecular disassembly, we concentrate on how the fragments connect with each other. RDKit  [^2] https://www.rdkit.org is an open-source Python software package providing chemical information. We use its built-in functions to determine the binding sites and connection modes between fragments.
@@
IUPAC Nomenclature in Organic Chemistry.   The international union of pure and applied chemistry (IUPAC)  (IUPAC, 1992) nomenclature system provides standardized naming conventions for chemical compounds, which is crucial for avoiding ambiguities arising from common or trivial names. This is particularly crucial in organic chemistry, where compound complexity can make identification challenging. When optimizing molecules guided by textual descriptions, we use the IUPAC names instead of SMILES strings to represent fragments, simultaneously complementing IUPAC information of the source molecules as molecular semantic representations. As shown in Fig.  3, on the one hand, IUPAC names provide a detailed and hierarchical description of a molecule’s structure, including functional groups, stereochemistry, branching, and chain length. This level of granularity can potentially enhance the precision of the generative process, especially when targeting specific chemical properties or structures. On the other hand, IUPAC nomenclature conventions provide more semantic information compared to SMILES. This is because they are more structured and descriptive than SMILES strings which feature a linear text format merely based on connectivity. The rich semantic details allow IUPAC names to convey more precise chemical information, making them more interpretable and suitable for language models that benefit from understanding chemical semantics.
@@
Textual Description Generation.   In addition to physical structure information, we consider the desired optimized chemical properties. Specifically, the original numerical property values and the changes between MMPs are supplemented into one textual description. Utilizing the above tools, we generate a textual description saturated with detailed molecular structure and rich semantic information. For instance, the target molecule C#CCOC(=O)c1ccccc1O can be described as: The target molecule is optimized from ethyl 2-hydroxybenzoate comprising that ethanol and 2-hydroxybenzaldehyde links via a O C SINGLE bond. Its logd changes between (0.3, 0.5], its solubility is about 2.442, and its clint is less than 1.904.
@@
3.3. SMILES Tokenizer
Inspired by Gong et al.  (Gong et al., 2024) who considered the fact that tokenizing every single a SMILES string character destroys the uniformity between multi-character units, we employ the same strategy to retain the semantic groups in SMILES strings. For instance, [CH3-] is a negatively charged methyl group rather than a sequence assembled by unrelated characters, such as [, C, and H. With all the semantic groups together forming our vocabulary, TransDLM is capable of encoding any SMILES string semantically and integrally, like CC([Si]=O)C[CH3-] being encoded as [[SOS],C,C,(,[Si],=,O,),C,[CH3-],[EOS],[PAD],…,[PAD]]. And we pad each sequence to the maximum length $n$.
@@ 3.4. Text-Guided Molecular Optimization
Unlike traditional diffusion models that sample from pure noise  (Li et al., 2022; Rombach et al., 2022), we propose a novel generation strategy for TransDLM, which samples from an encoded source molecule SMILES string, as mentioned in Chapter  3.3. Our approach effectively addresses the uncertainty and infeasibility of previous MO methods and adjusts the orientation to the desired optimized molecules.
For the purpose of converting textual description into machine-readable language, a pre-trained language model is utilized to map the text sequence $C$ to its latent embeddings $\mathbf{C}\in\mathbb{R}^{d_{1}\times m}$, where $d_{1}$ represents the embedding dimension of the language model output. Our model is based on transformer framework, which functions as $f_{\theta}$ in Eq. ( 2). Considering the current diffusion state $x_{t}$, time step $t$, and text embeddings $\mathbf{C}$, the initial state $x_{0}$ is predicted as $\hat{x_{0}}=f_{\theta}(x_{t},t,\mathbf{C})$. Given that self-attention mechanism is disordered, the transformer architecture does not inherently have the location information. Hence, the transformer loses the input’s sequential information. As a result, we adopt a position encoding strategy  (Vaswani, 2017) in the first layer $z_{t}^{(0)}$ to enable our model to recognize the position of each element in the input sequence, defined as:
@@
where $PosEmb$ is the positional embedding, $TimeEmb$ is an operation to embed the time step $t$ into a high-dimensional vector, $W_{in}\in\mathbb{R}^{d_{2}\times d}$, $z_{t}^{(0)}\in\mathbb{R}^{d_{2}\times n}$, and $d_{2}$ is the dimension for transformer.
In the standard transformer self-attention mechanism, each token performs attention calculations on others in the same input sequence. Therefore, to improve the multi-modal data processing capability of our model and enhance the positive guidance of textual description to MO, TransDLM uses the cross-attention mechanism to calculate the input sequence with textual description as an additional context. Specifically, our transformer backbone is composed of $L$ layers with a cross-attention block respectively, which incorporates the textual description into the hidden states as follows:
@@
where $W_{*}^{(i)}\in\mathbb{R}^{d_{2}\times d_{2}}$ represents the learnable parameters in the $i$-th layer, and $MLP$ is the multilayer perception.
After completing the above procedures, we obtain the matrix $\hat{x}_{0}$ from the initial embedded source molecule, guided by the textual description $C$. Then, the matrix $\hat{x}_{0}$ is prepared for rounding and subsequently transformed into a SMILES string.
Algorithm 1 Training algorithm
  repeat
     Sample $M$, $\mathbf{C}$ and $t$
     Acquire $x_{0}$ through embedding process
     if $t=0$ then
        $g=\nabla_{\theta}(-\log{p_{\theta}(M|x_{0})})$
     else
        $g=\nabla_{\theta}\|f_{\theta}(x_{t},t,\mathbf{C})-x_{0}\|^{2}$
     end if
     Implement optimization through gradient $g$
  until converged
@@ 3.5. TransDLM Training
During the training process, our methodology mainly maximizes the variational lower bound (VLB) of the marginal likelihood. Moreover, we refine and adapt our approach according to the insights from previous studies (Ho et al., 2020; Li et al., 2022). As the training procedure shown in Algorithm  1, we primarily aim to train the neural network, denoted as $f_{\theta}$, to progressively reconstruct the original data $x_{0}$ at every time step within the denoising diffusion trajectory, described as:
(5) $\mathcal{L}_{M,C}=\mathop{\mathbb{E}}\limits_{q(x_{0:T}|M)}\left[\sum_{t=1}^{T}\|f_{\theta}(x_{t},t,\mathbf{C})-x_{0}\|^{2}-\log{p_{\theta}(M|x_{0})}\right],$
where $p_{\theta}(M|x_{0})$ represents the rounding process, featuring a product of the softmax distribution $p_{\theta}(M|x_{0})=\prod_{i=0}^{n}p(a_{i}|x_{0[:,i]})$.
By iteratively refining the noisy intermediate states to their original forms, the TransDLM learns to effectively reverse the diffusion process, leveraging underlying data structure and guiding text.
@@ 4. THEORETICAL ANALYSIS
4.1. Gradient Error Calculation
The loss function of traditional methods usually consists of attribute and structure loss, which can be expressed as:
(6) $L_{\text{trad}}(M)$ $=\alpha\cdot L_{\text{property}}(M)+\beta\cdot L_{\text{structure}}(M)$
$=\alpha\cdot(\hat{y}(M)-y_{\text{target}})^{2}+\beta\cdot\|M-M_{\text{target}}\|^{2},$
where $\hat{y}(M)$ is the predicted attribute value of the molecule $M$ by the external property predictor, $M_{\text{target}}$ denotes the target molecular structure, and $\alpha$ and $\beta$ are hyperparameters that adjust the attribute and structure losses’ weights. As a result, inferred from Eqs. ( 5) and ( 6), the respective gradient of molecule $M$ are:
(7) $\nabla_{\theta}\mathcal{L}_{M,C}=$ $2\mathbb{E}_{q(x_{0:T}|M)}\left[(f_{\theta}(x_{t},t,\mathbf{C})-x_{0})\nabla_{\theta}f_{\theta}(x_{t},t,\mathbf{C})\right]$
$\qquad\qquad\qquad\qquad\qquad\,-\nabla_{\theta}\log p_{\theta}(M|x_{0}),$
(8) $\nabla_{M}L_{\text{trad}}$ $=\alpha\cdot\nabla_{M}L_{\text{property}}+\beta\cdot\nabla_{M}L_{\text{structure}}$
$=2\alpha(\hat{y}(M)-y_{\text{target}})\nabla_{M}\hat{y}(M)+2\beta(M-M_{\text{target}}),$
Errors occur in the attribute loss’s gradient due to the ones within the external property predictor $\hat{y}(M)$. Thus, we define the following:
  • Attribute prediction error: $\epsilon_{\text{ext}}=\hat{y}(M)-y(M)$.
  • Attribute gradient prediction error: $\epsilon_{\text{grad}}=\nabla_{M}\hat{y}(M)-\nabla_{M}y(M)$.
  • Actual attribute values and gradients: $y(M)$ and $\nabla_{M}y(M)$.
Therefore, the error of attribute loss gradient is:
(9) $\epsilon_{\text{grad, property}}=\nabla_{M}L_{\text{property}}-\nabla_{M}L_{\text{property}}^{\text{true}}$
$=2\big{[}(\epsilon_{\text{ext}})\nabla_{M}y(M)+(y(M)-y_{\text{target}})\epsilon_{\text{grad}}+\epsilon_{\text{ext}}\epsilon_{\text{grad}}\big{]},$
The gradient error of our model mainly arises from the approximation error. Due to the direct fitting of $x_{0}$ and the use of text prompts $\mathbf{C}$ containing attribute information, our model’s gradient error is relatively small, which is defined as:
(10) $\epsilon_{\text{grad, our}}=2(f_{\theta}(x_{t},t,\mathbf{C})-x_{0})\nabla_{\theta}f_{\theta}(x_{t},t,\mathbf{C}),$
Finally, we compare our method’s gradient error variance with traditional approaches.
4.2. Comparison of Gradient Error Variance
For traditional methods, the attribute loss gradient error is:
(11) $\sigma_{\text{grad, trad}}^{2}$ $=4\alpha^{2}\Big{[}\operatorname{Var}\left(\epsilon_{\text{ext}}\nabla_{M}y(M)\right)$
$+\operatorname{Var}\left((y(M)-y_{\text{target}})\epsilon_{\text{grad}}\right)+\operatorname{Var}\left(\epsilon_{\text{ext}}\epsilon_{\text{grad}}\right)\Big{]}$
$=4\alpha^{2}\Big{[}(\nabla_{M}y(M))^{2}\sigma_{\text{ext}}^{2}+(y(M)-y_{\text{target}})^{2}\sigma_{\text{grad}}^{2}$
$\qquad\qquad\qquad\qquad\quad\quad\quad\quad\quad\ \ +\sigma_{\text{ext}}^{2}\sigma_{\text{grad}}^{2}\Big{]},$
where we assume $\epsilon_{\text{ext}}$ and $\epsilon_{\text{grad}}$ are independent and have an expected value of zero and $y(M)-y_{\text{target}}$ and $\nabla_{M}y(M)$ are deterministic or their variances can be ignored.
Similarly, our TransDLM’s gradient error variance is:
(12) $\sigma_{\text{grad, our}}^{2}$ $=4\operatorname{Var}\left((f_{\theta}(x_{t},t,\mathbf{C})-x_{0})\nabla_{\theta}f_{\theta}(x_{t},t,\mathbf{C})\right)$
$=4\Big{[}(\mathbb{E}[\nabla_{\theta}f_{\theta}(x_{t},t,\mathbf{C})])^{2}\sigma_{\text{model}}^{2}$
$\qquad\qquad\ +\sigma_{\text{model}}^{2}\sigma_{\text{grad, model}}^{2}\Big{]},$
For traditional methods, the gradient error variance depends on the model’s gradient calculation error $\sigma_{\text{grad}}^{2}$ and the error $\sigma_{\text{ext}}^{2}$ of the external property predictors. Due to the potential for significant errors in external predictors, this can significantly increase the overall gradient error variance. On the contrary, the endogenous errors in TransDLM are mainly controlled by $\sigma_{\text{model}}^{2}$ and $\sigma_{\text{grad, model}}^{2}$, and do not rely on external property predictors. As a consequence, the variance term caused by external errors is reduced.
Assuming that the fitting ability of traditional models and TransDLM is the same, that is, $(y(M)-y_{\text{target}})^{2}$ and $\mathbb{E}[\nabla_{\theta}f_{\theta}(x_{t},t,\mathbf{C})]$ are similar, conventional methods have additional error terms related to external property predictors. We can infer that $(\nabla_{M}y(M))^{2}\sigma_{\text{ext}}^{2}$ and $\sigma_{\text{ext}}^{2}\sigma_{\text{grad}}^{2}$ in $\sigma_{\text{grad, trad}}^{2}$ are additional error terms excluded from diffusion models. Therefore, as long as $\sigma_{\text{ext}}^{2}$ is not a very small value, $\sigma_{\text{grad, trad}}^{2}$ will be larger than $\sigma_{\text{grad, our}}^{2}$. All reasonings above taken into account, our text-guided MO model with the diffusion language model outperforms traditional methods in terms of controlling error propagation, that is $\sigma_{\text{grad, our}}^{2}\textless\sigma_{\text{grad, trad}}^{2}$.
@@ 5.1. Dataset
Based on our research purpose, our evaluation centers on an MMP dataset  (He et al., 2021), which is currently the sole publicly available dataset. This dataset includes 198,558 source-target molecule pairs with their ADMET properties, and we ensured consistent dataset partitioning for all the experiments.
Table 1. Optimization results for structural similarity compared with baseline models and ablation studies on the test split of MMP dataset. Bold indicates the best scores and underlined signifies the second best scores. Additionally, $\mathrm{TransDLM}_{\mathrm{noise}}$ and $\mathrm{TransDLM}_{\mathrm{SMILES}}$ respectively denote sampling from pure noise and optimizing guided by SMILES string-based descriptions.
@@
TransDLM 0.74 0.009 14.838 0.818 0.792 0.665 0.109 0.993
Table 2. Optimization results for ADMET properties compared with baseline models and ablation studies on the test split of MMP dataset. The criterion ’All’ indicates the ratio of optimized molecules satisfying all property requirements.
@@
MIMOSA 0.1 0.765 0.755 0.075
Modof 0.08 0.615 0.733 0.044
MolSearch 0.06 0.706 0.746 0.039
Chemformer 0.12 0.769 0.755 0.086
FRATTVAE 0.044 0.663 0.745 0.026
TransDLM 0.157 0.783 0.757 0.11
$\mathrm{TransDLM}_{\mathrm{noise}}$ 0.16 0.789 0.757 0.108
$\mathrm{TransDLM}_{\mathrm{SMILES}}$ 0.038 0.756 0.753 0.025

### v3

@@ 3.2. Description generator
Molecular structure feature.   Molecular structure feature effectively enhances insights into molecules and model performance in specific tasks (Ferreira et al., 2015; Wang et al., 2024b; Djeddi et al., 2023). Among them, fragment-based molecular structure feature has garnered substantial attention from researchers (Green et al., 2021; Yang et al., 2021). Specifically, we disassemble molecules using RECAP algorithm (Lewell et al., 1998) to obtain a series of chemically reasonable fragments (for more details in Appendix B). After molecular disassembly, we utilize RDKit [^2] https://www.rdkit.org to analyze which atoms on the fragments are involved in the connection and the types of chemical bonds formed between fragments.
@@
IUPAC nomenclature in organic chemistry.   The international union of pure and applied chemistry (IUPAC) (IUPAC, 1992) nomenclature system provides standardized naming conventions for chemical compounds, which is crucial for avoiding ambiguities arising from common or trivial names. When optimizing molecules guided by textual descriptions, we represent fragments using IUPAC names instead of SMILES, and also use the IUPAC information of the source molecules as semantic representations. As shown in Fig. 3, IUPAC names offer a hierarchical description of molecular structure, covering functional groups, stereochemistry, branching, and chain length. This finer granularity can improve the precision of generation, especially when targeting specific chemical properties or structures. Compared with the linear, connectivity-based SMILES format, IUPAC nomenclature carries clearer semantic information and is more interpretable for language models. For example, the SMILES C(=O)O does not explicitly indicate a carboxyl group to a general language model. In contrast, the descriptive nature of IUPAC names allows models to better capture chemical meaning. Notably, we employ an external tool [^3] https://www.chemtools.cn/iupac/smilesToIupac.HTML to convert SMILES strings into IUPAC names.
@@
Textual description generation.   In addition to physical structure information, we consider the desired optimized chemical properties. Specifically, the original numerical property values and the changes between MMPs are supplemented into one textual description. Utilizing the above tools, we generate a textual description saturated with detailed molecular structure and rich semantic information. The grammatical template for textual description is as below:
The target molecule is optimized from [IUPAC name of source molecule] comprising that [IUPAC name of fragment 1] and [IUPAC name of fragment 2] links via a [bond type] bond, $\cdots$. Its [property 1] [change verb + value/range], [property 2] [change verb + value/range] $\cdots$.
For instance, the target molecule C#CCOC(=O)c1ccccc1O can be described as: The target molecule is optimized from ethyl 2-hydroxybenzoate comprising that ethanol and 2-hydroxybenzaldehyde links via a O C SINGLE bond. Its logd changes between (0.3, 0.5], its solubility is about 2.442, and its clint is less than 1.904.
3.3. SMILES tokenizer
Inspired by Gong et al. (Gong et al., 2024) who considered that tokenizing every single character of a SMILES string destroys the uniformity between multi-character units, we employ the same strategy to retain the semantic groups in SMILES strings. For instance, [CH3-] is a negatively charged methyl group rather than a sequence assembled by unrelated characters, such as [, C, and H. With all the semantic groups together forming our vocabulary, TransDLM is capable of encoding any SMILES string semantically and integrally, like CC([Si]=O)C[CH3-] being encoded as [[SOS],C,C,(,[Si],=,O,),C,[CH3-],[EOS],[PAD], …,[PAD]]. And we pad each sequence to $n$.
@@ 3.4. Text-guided molecular optimization
Unlike traditional diffusion models that sample from pure noise (Li et al., 2022; Rombach et al., 2022), TransDLM generates molecules by sampling from encoded source molecule SMILES strings. This strategy mitigates the uncertainty and infeasibility of previous MO methods and better guides generation toward the desired target.
Our model is based on a transformer backbone, which functions as $f_{\theta}(\cdot)$ in Eq. 3 and predicts $\hat{x}_{0}=f_{\theta}(x_{t},t,\mathbf{C})$. Since self-attention mechanism is permutation-invariant, the model lacks positional awareness. To preserve sequence order, we apply positional and temporal embeddings (Vaswani, 2017) in the first layer $z_{t}^{(0)}$, defined as:
@@
where $PosEmb$ denotes positional encoding, $TimeEmb(\cdot)$ embeds diffusion step $t$, and $W_{in}\in\mathbb{R}^{d_{2}\times d}$ projects the input to transformer dimension $d_{2}$.
In the standard transformer self-attention mechanism, each token performs attention calculations on others in the same input sequence. Therefore, to improve the multi-modal data processing capability of our model and enhance the positive guidance of textual description to MO, TransDLM incorporates cross-attention into each of the $L$ transformer layers. This enables the latent molecule representation to attend to the description embedding $\mathbf{C}$:
@@
where $W_{*}^{(i)}\in\mathbb{R}^{d_{2}\times d_{2}}$ represents the learnable parameters in the $i$-th layer, and $MLP(\cdot)$ is a multilayer perception.
After completing the above procedures, we obtain the matrix $\hat{x}_{0}$ from the initial embedded source molecule, guided by the textual description $C$. Then, $\hat{x}_{0}$ is rounded to produce the final SMILES string.
@@ 3.5. TransDLM training
During the training process, our methodology mainly maximizes the variational lower bound (VLB) of the marginal likelihood. Moreover, we refine and adapt our approach according to the insights from previous studies (Ho et al., 2020; Li et al., 2022). We primarily aim to train $f_{\theta}(\cdot)$ to progressively reconstruct the desired data $x_{0}$ at every diffusion step within the denoising diffusion trajectory, described as:
(6) $\mathcal{L}_{M,C}=\mathop{\mathbb{E}}\limits_{q(x_{0:T}\mid M)}\left[\sum_{t=1}^{T}\|f_{\theta}(x_{t},t,\mathbf{C})-x_{0}\|^{2}-\log{p_{\theta}(M\mid x_{0})}\right],$
where $p_{\theta}(M\mid x_{0})$ represents the rounding process, featuring a product of the softmax distribution $p_{\theta}(M\mid x_{0})=\prod_{i=0}^{n}p(a_{i}\mid x_{0[:,i]})$.
By iteratively refining the noisy intermediate states to optimized forms, the TransDLM learns to effectively reverse the diffusion process, leveraging underlying data structure and guiding text.
@@ 4. Theoretical analysis
4.1. Error propagation in traditional methods
Usually, the optimization process of traditional methods involves $Z$ iterations. The estimated property values at iteration $z$ using an external predictor is $\hat{y}_{z}=y_{z}+\epsilon_{z}$, where $y_{z}$ represents the true property value, and $\epsilon_{z}\sim\mathcal{N}(0,\sigma^{2})$ is the prediction error. Since the optimization direction is determined by the gradient $\nabla_{\theta}\mathcal{L}(\hat{y}_{z})$, the parameters are updated to $\theta_{z+1}=\theta_{z}-\eta\nabla_{\theta}\mathcal{L}(\hat{y}_{z})$, where $\eta$ represents the learning rate. Consequently, the projection of error in the gradient direction is:
(7) $\Delta_{z}=\nabla_{\theta}\mathcal{L}(y_{z})-\nabla_{\theta}\mathcal{L}(\hat{y}_{z})\approx\epsilon_{z}\cdot\nabla_{\theta}^{2}\mathcal{L}(\hat{y}_{z}).$
Assuming the upper bound of the spectral radius of the Hessian matrix is $H$, the cumulative error can be expressed as follows:
(8) $\mathcal{E}_{\text{trad}}\leq\sum_{z=1}^{Z}\eta H|\epsilon_{z}|\sim O(Z\eta H\sigma),$
which implies that $\mathcal{E}_{trad}$ will increase linearly with the number of iterations $Z$, leading to a substantial deviation of the final optimization direction from the theoretically optimal value.
4.2. Error suppression mechanism of TransDLM
On the contrary, TransDLM exhibits exceptional error suppression capabilities. The generation distribution at each diffusion step $t$ is given by $\mathcal{N}(\mu_{\theta}(x_{t},t,\mathbf{C}),\sum_{t})$, where $\mu_{\theta}$ is influenced by condition $\mathbf{C}$, and $\sum_{t}$ represents a fixed value. The error associated with TransDLM at time $t$ arises from the deviation in predicting $x_{0}$, expressed as $\delta_{t}=\hat{x}_{0}^{(t)}-x_{0}$, where $\hat{x}_{0}^{(t)}=f_{\theta}(x_{t},t,\mathbf{C})$. Consequently, the discrepancy between the actual generated value $x_{t-1}$ and its ideal counterpart $x_{t-1}^{*}$ is:
(9) $\Delta x_{t-1}=\mu_{\theta}-\mu^{*}=\frac{\sqrt{\overline{\alpha}_{t-1}}\beta_{t}}{1-\overline{\alpha}_{t}}\delta_{t}.$
Notably, as outlined in the diffusion scheduling design (Ho et al., 2020), the coefficient $\gamma_{t}=\frac{\sqrt{\overline{\alpha}_{t-1}}\beta_{t}}{1-\overline{\alpha}_{t}}<1$, indicating that the propagation of error $\delta_{t}$ is significantly attenuated.
Due to the Markovian nature of the diffusion process (Cao et al., 2024), $x_{t-1}$ is solely dependent on $x_{t}$ and the currently predicted value of $x_{0}$. The influence of errors on subsequent steps is expressed as follows:
(10) $\Delta x_{t-1}$ $\propto\gamma_{t}\delta_{t}$
$\Delta x_{t-2}$ $\propto\gamma_{t}\gamma_{t-1}\delta_{t}+\gamma_{t-1}\delta_{t-1}$
$\cdots$
As a consequence, the cumulative error of TransDLM is:
(11) $\mathcal{E}_{\text{ours}}\propto\sum_{t=1}^{T}\Big(\prod_{s=1}^{t}\gamma_{s}\Big)\delta_{t}.$
Due to the exponential decay of $\prod_{s=1}^{t}\gamma_{s}$ with respect to $t$, error propagation is effectively suppressed.
4.3. Comparative analysis
For traditional methods, error propagation is characterized as a random walk process. On the contrary, in the TransDLM framework, errors not only accumulate independently due to the fixed covariance matrix $\sum_{t}$, but also the series presented in Eq. 11 converge owing to the condition $\gamma_{s}<1$. Assuming the variance of single-step errors for both methods is $\sigma^{2}$, we can readily derive the total error variance for each approach:
(12) $Var(\theta_{Z})=\sum_{z=1}^{Z}Var(\Delta_{z})=Z^{2}\eta^{2}H^{2}\sigma^{2}\sim O(Z^{2}\sigma^{2}),$
(13) $Var(x_{0})=\sum_{t=1}^{T}\Big(\prod_{s=1}^{t}\gamma_{s}\Big)\sigma^{2}\leq T\sigma^{2}\sim O(T\sigma^{2}).$
For traditional methods relying on external property predictors, the total variance exhibits a quadratic growth pattern with respect to the number of optimization steps $Z$. This implies that as the optimization process is extended, error accumulation becomes more pronounced compared to TransDLM, resulting in substantial deviations from the optimal solution. In contrast, TransDLM benefits from an inherent error suppression mechanism within its diffusion framework, which ensures that accumulated errors remain significantly lower than those observed in traditional approaches. The derivation presented above offers robust theoretical evidence that TransDLM effectively mitigates error propagation, thereby ensuring a more stable and reliable multi-property MO process.
@@ 5.1. Dataset
Based on our research purpose, our evaluation centers on an MMP dataset (He et al., 2021), which is currently the sole publicly available dataset and includes 198558 source–target molecule pairs with their ADMET properties. We randomly split the whole dataset into 90 % as training and validation, and 10 % as test, and further split the former into 90 % as training and 10 % as validation, with ensuring that the random seed and dataset partitioning for all experiments is consistent.
Table 1. Optimization results for structural similarity compared with baseline models on the test split of MMP dataset. Bold indicates the best scores.
@@
DyMol 0.369 0 35.728 0.406 0.365 0.254 5.155 1
TransDLM (Ours) 0.74 0.009 14.838 0.818 0.792 0.665 0.109 0.993
Table 2. Optimization results for ADMET properties compared with baseline models on the test split of MMP dataset. The criterion ’All’ indicates the ratio of optimized molecules satisfying all property requirements.
@@
MIMOSA 0.1 0.765 0.755 0.075 0.911 0.141
Modof 0.08 0.615 0.733 0.044 0.908 0.145
MolSearch 0.06 0.706 0.746 0.039 0.923 0.136
FRATTVAE 0.044 0.663 0.745 0.026 0.865 0.214
DyMol 0.041 0.58 0.722 0.023 0.876 0.207
TransDLM (Ours) 0.157 0.783 0.757 0.11 0.929 0.121